OpenStudy (anonymous):

HELPPPPPP PLZZZZ!!! in a right triangle, find the length of the side not given.. give an exact answer and an approximation to three decimal places. a=15, b=15, c=??? what is the exact value of c? type an exact answer using radicals as needed.

OpenStudy (anonymous):

use a^2 + b^2 = c^2

OpenStudy (anonymous):

15 ^2 = 225 225 +225 = 450 c^2 = 450 c = \[\sqrt{450}\

OpenStudy (anonymous):

Using Pythagorean's Theorem, find c. \[a^{2}+b^{2}=c^{2}\] \[15^{2}+15^{2}=c^{2}\] \[450=c^{2}\] \[\sqrt{450}=c\]

OpenStudy (anonymous):

Ninja'd =P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

then it asks what is the value of c approximated to 3 decimal places?

OpenStudy (anonymous):

21.213

myininaya (myininaya):

enter it into your calculator round to nearest thousandths

OpenStudy (anonymous):

Get a calculator to do that (I don't have one handy)

OpenStudy (anonymous):

haha im doing my math hmrk :P and im stuck with advanced math question? help?

OpenStudy (anonymous):

Post it, I'll try.

OpenStudy (anonymous):

thank u everyone!! im ganna post another one!! i only have 4 left!!